### v2

# Scaling laws for velocity profile of granular flow in rotating drums

Abstract
We numerically analyze the steady flow of two-dimensional granular materials in a rotating drum using the discrete element method and a continuum model. The velocity fields obtained from both methods are quantitatively consistent. Two distinct regions exist in the granular flow: the surface flow layer and the static flow regime associated with the rigid rotation near the bottom. The thickness of the surface flow layer increases with the drum diameter and shows a slight dependence on the angular velocity of the rotating drum. We derived scaling laws for the velocity profile and surface layer thickness using dimensional analysis applied to the continuum equations. The validity of the scaling laws is confirmed numerically.
I Introduction
Flows in rotating drums are used in various industrial processes of granular materials, including mixing, breaking, granulating, drying, and chemical reactions. The form of the flow inside the drum depends on the Froude number $Fr=\Omega^{2}D/2g$ [1], where $\Omega$ is the angular velocity of the rotation, $D$ is the diameter of the drum, and $g$ is the gravitational acceleration. For $10^{-4}<Fr<10^{-2}$, a steady flow driven by the rotating drum occurs, as shown in Fig. 1. In this steady state, the flow field comprises two regions: the “surface flow layer” and the “static flow regime.” In the surface flow layer, the granular material flows down along the slope of the free surface. In the static flow regime, the granular material exhibits a solid-like rigid rotation following the drum wall.
Numerous experimental studies have investigated the dependence of the surface flow layer thickness $h$ on $D$ and $\Omega$ in drums half-filled with granular materials [2, 3, 4, 5, 6]. In. Ref. [2], it is reported that $h$ is proportional to $D$ and almost independent of $\Omega$ in a drum long enough in the axial direction. According to Ref. [5], $h$ in axially thin rotating drums exhibits a power-law dependence on $\Omega$ for sufficiently large $D/d$ with the particle diameter $d$. The experiments in Ref. [7] suggested that $h$ in thin drums obeys $h/d\propto\tilde{Q}^{0.44}$ with the dimensionless flow rate $\tilde{Q}=\frac{1}{2}Fr^{1/2}\left(D/d\right)^{3/2}$. However, the dependence of $h$ has not been explained theoretically.
[…]
Figure 1: Schematic of the steady granular flow in a rotating drum.
Numerical simulations have been conducted to analyze the granular flow in rotating drums. Several researchers have used the discrete element method (DEM) [8, 9, 10, 11, 12], where the equation of motion for each particle is solved numerically. However, the numerical cost for the DEM simulation increases with the number of particles in the system with $D\gg d$. Therefore, direct simulations of the systems used in industrial applications with a large number of particles are challenging. In recent years, simulation methods based on continuum equations have been proposed, which are free from the restriction on the number of particles. In Refs. [13, 14, 15], granular materials are treated as elastoplastic bodies to investigate their mixing driven by rotating drums. In Refs. [16, 17], the computational fluid dynamics (CFD) method was used to study the flows in rotating drums, where granular materials are treated as non-Newtonian viscous fluids. It has been reported that the velocity field in the rotating drums is reproduced by these continuum approaches; however, its dependence on $D$ and $\Omega$ has not been studied in detail.
Various constitutive relations have been utilized in these continuum approaches. Among the constitutive relations, the $\mu(I)$-rheology has been applied to various granular flows because of its simplicity [18, 19]. In the $\mu(I)$-rheology, the shear stress $\tau$ of granular materials under uniform shear with the shear rate $\dot{\gamma}$ and pressure $p$ is given by
$\tau=\mu(I)p.$ (1)
Here, $\mu(I)$ is the bulk friction depending on the inertial number $I=\dot{\gamma}/\sqrt{p/\rho_{s}d^{2}}$, where $\rho_{s}$ is the density of the particles comprising the granular materials. In Ref. [19], a functional form of $\mu(I)$ is proposed as follows:
$\mu\left(I\right)=\mu_{s}+\frac{\mu_{2}-\mu_{s}}{I_{0}/I+1},$ (2)
where $\mu_{s}$, $\mu_{2}$, and $I_{0}$ are constants depending on the elasticity, dissipation, and friction of the particles [18]. The continuum model using the $\mu(I)$-rheology treats granular materials as non-Newtonian fluids with the viscosity $\eta=\mu(I)p/\dot{\gamma}$. Recently, the constitutive relations incorporating non-local effects in the $\mu(I)$-rheology have been proposed [20, 21, 22]. The continuum models using the $\mu(I)$-rheology reproduce various granular flows, including those in rotating drums [23, 13, 14, 16]. Therefore, this constitutive law is expected to explain the dependence of the velocity profile on $D$ and $\Omega$.
This study aims to understand the steady flow of granular materials in rotating drums. We focus on 2D rotating drums half-filled with granular materials for $D/d\gg 1$ and obtain the velocity field numerically by DEM and CFD simulations. The details of the numerical simulations are described in Sec. II. The dependence of the velocity field on $D$ and $\Omega$ is shown in Sec. III, where we demonstrate that the continuum model quantitatively reproduces the flow obtained in the DEM simulation over a broad parameter space. In Sec. IV, scaling laws characterizing the velocity field are derived analytically from the continuum model using the $\mu(I)$-rheology and verified numerically. The method for calculating the velocity field in the DEM simulation is presented in Appendix A. The values of the parameters $\mu_{s}$, $\mu_{2}$, and $I_{0}$ for the $\mu(I)$-rheology are estimated in Appendix B. The packing fraction of granular materials in rotating drums is demonstrated in Appendix C. The position of the free surface is discussed in Appendix D. The details of the derivation for the analytical expression of $h$ are shown in Appendix E.
II Numerical Method
II.1 Setup
In this paper, we consider a two-dimensional granular flow in a rotating drum, as illustrated in Fig. 1. The origin of the Cartesian coordinate system $(X,Z)$ is placed at the center of the drum with diameter $D$. The polydisperse particles with the minimum diameter $d$ are half-filled in the drum. The drum is rotated counterclockwise with the angular velocity $\Omega$. The drum diameter $D$ and angular velocity $\Omega$ are set to $D=150d$, $300d$, and $450d$ and $1.26\times 10^{-3}\leq\Omega/\sqrt{g/d}\leq 1.26\times 10^{-2}$ to satisfy $10^{-4}<Fr<10^{-2}$, respectively.
II.2 DEM simulation
II.2.1 Time evolution equations
In the DEM simulation [24, 25, 26], we consider $N$ particles in a drum. The dynamics of particle $i$ is given by the time evolution equations for the position ${\bf\it r}_{i}=(X_{i},Z_{i})$ and angular velocity $\omega_{i}$ of particle $i$ as follows:
$m_{i}\frac{d^{2}{\bf\it r}_{i}}{dt^{2}}$ $=\sum_{j\neq i}{\bf\it F}_{ij}+m_{i}{\bf\it g},$ (3)
$I^{\rm p}_{i}\frac{d\omega_{i}}{dt}$ $=\sum_{j\neq i}T_{ij}$ (4)
with ${\bf\it g}=(0,-g)$. Here, $m_{i}$ and $I^{\rm p}_{i}$ are the mass and moment of inertia of particle $i$, respectively. The contact force and torque between particle $i$ and particle $j$ are denoted by ${\bf\it F}_{ij}$ and $T_{ij}$, respectively.
The contact force ${\bf\it F}_{ij}$ consists of the normal component ${\bf\it F}^{(\mathrm{n})}_{ij}$ and the tangential component ${\bf\it F}^{(\mathrm{t})}_{ij}$ as follows:
${\bf\it F}_{ij}={\bf\it F}^{(\mathrm{n})}_{ij}+{\bf\it F}^{(\mathrm{t})}_{ij}.$ (5)
The normal contact force ${\bf\it F}^{(\mathrm{n})}_{ij}$ is given by
${\bf\it F}^{(\mathrm{n})}_{ij}=\left(-k^{(\mathrm{n})}{\bf\it\delta}^{(\mathrm{n})}_{ij}-\eta^{(\mathrm{n})}_{ij}{\bf\it v}^{(\mathrm{n})}_{ij}\right)\Theta\left(\delta_{ij}\right),$ (6)
where $k^{(\mathrm{n})}$ and $\eta^{(\mathrm{n})}_{ij}$ are the normal elastic and the normal viscous constants, respectively; here, $\Theta(x)$ is the Heaviside step function satisfying by $\Theta(x)=1$ for $x\geq 0$, and $\Theta(x)=0$ for otherwise. The normal displacement ${\bf\it\delta}^{(\mathrm{n})}_{ij}$, the normal relative velocity ${\bf\it v}^{(\mathrm{n})}_{ij}$, and the normal unit vector ${\bf\it n}_{ij}$ in Eq. (6) are respectively given by
${\bf\it\delta}^{(\mathrm{n})}_{ij}=\delta_{ij}{\bf\it n}_{ij},$ (7)
${\bf\it v}^{(\mathrm{n})}_{ij}=\left\{({\bf\it v}_{i}-{\bf\it v}_{j})\cdot{\bf\it n}_{ij}\right\}{\bf\it n}_{ij},$ (8)
${\bf\it n}_{ij}=\frac{{\bf\it r}_{j}-{\bf\it r}_{i}}{\left|{\bf\it r}_{j}-{\bf\it r}_{i}\right|},$ (9)
where ${\bf\it v}_{i}$ is the velocity of particle $i$, and $\delta_{ij}$ is the overlap between particles $i$ and $j$ given by
$\delta_{ij}=\frac{d_{i}+d_{j}}{2}-\left|{\bf\it r}_{j}-{\bf\it r}_{i}\right|$ (10)
with the diameter $d_{i}$ of particle $i$.
The tangential contact force ${\bf\it F}^{(\mathrm{t})}_{ij}$ obeys the Coulomb friction law as follows:
${\bf\it F}^{(\mathrm{t})}_{ij}=\mathrm{min}\left(\left|{\bf\it F}^{(\mathrm{t,st})}_{ij}\right|,\;\mu_{\mathrm{p}}\left|{\bf\it F}^{(\mathrm{n})}_{ij}\right|\right)\frac{{\bf\it F}^{(\mathrm{t,st})}_{ij}}{\left|{\bf\it F}^{(\mathrm{t,st})}_{ij}\right|}\Theta\left(\delta_{ij}\right)$ (11)
with the particle friction coefficient $\mu_{\mathrm{p}}$. Here, $\mathrm{min}(a,b)$ in Eq. (11) selects the smaller of $a$ and $b$, and ${\bf\it F}^{(\mathrm{t,st})}_{ij}$ is given by
${\bf\it F}^{(\mathrm{t,st})}_{ij}=-k^{(\mathrm{t})}\delta^{(\mathrm{t})}_{ij}{\bf\it t}_{ij}-\eta^{(\mathrm{t})}_{ij}{\bf\it v}^{(\mathrm{t})}_{ij}$ (12)
with the tangential spring constant $k^{(\mathrm{t})}$ and the tangential viscous constant $\eta^{(\mathrm{t})}_{ij}$. The tangential displacement ${\delta}^{(\mathrm{t})}_{ij}$, the tangential relative velocity ${\bf\it v}^{(\mathrm{t})}_{ij}$, and the tangential unit vector ${\bf\it t}_{ij}$ are given by
[…]
${\bf\it v}^{(\mathrm{t})}_{ij}=\left\{\left({\bf\it v}_{i}-{\bf\it v}_{j}\right)\cdot{\bf\it t}_{ij}+\frac{1}{2}\left(d_{i}\omega_{i}+d_{j}\omega_{j}\right)\right\}{\bf\it t}_{ij},$ (14)
${\bf\it t}_{ij}=\left(-n_{ij,Z},\;n_{ij,X}\right).$ (15)
The subscript “$\mathrm{stick}$” in Eq. (13) indicates that the integral is performed only when $\left|{\bf\it F}^{(\mathrm{t,st})}_{ij}\right|<\mu_{\mathrm{p}}\left|{\bf\it F}^{(\mathrm{n})}_{ij}\right|$. The torque $T_{ij}$ is given by
$T_{ij}=\frac{d_{i}}{2}{\bf\it F}^{(\mathrm{t})}_{ij}\cdot{\bf\it t}_{ij}.$ (16)
II.2.2 Numerical conditions
The diameters of the particles range uniformly from $d$ to $1.2d$ [10]. The mass and moment of inertia of particle $i$ are given by $m_{i}=\frac{1}{4}\pi\rho_{s}d_{i}^{2}$ and $I^{p}_{i}=\frac{1}{8}m_{i}{d_{i}}^{2}$, respectively. We set $k^{(\mathrm{n})}=2.0\times 10^{5}mg/d$, $k^{(\mathrm{t})}/k^{(\mathrm{n})}=2/7$, and $\mu_{p}=0.50$, where $m$ is the mass of the particle with diameter $d$ [27]. The normal viscous constant between particles $i$ and $j$ is given by
$\eta^{(\mathrm{n})}_{ij}=\eta^{(\mathrm{t})}_{ij}=-2\ln{e}\sqrt{\frac{m^{*}_{ij}k^{(\mathrm{n})}}{\pi^{2}+(\ln{e})^{2}}}$ (17)
with the coefficient of restitution $e=0.92$ and
$m^{*}_{ij}=\frac{m_{i}m_{j}}{m_{i}+m_{j}}.$ (18)
The time evolution equations are numerically solved using the leapfrog algorithm with the time step $\Delta t=1.0\times 10^{-4}\sqrt{d/g}$.
The rough wall surface of the drum is modeled by placing $N_{\rm w}$ particles of diameter $d$ equally spaced on the circle of diameter $D$ to avoid slip at the surface. The particles constituting the rough wall are rotated around the circle with the angular velocity $\Omega$. We set $N_{\rm w}=475$, $632$, and $946$ for the drums with $D=150d,300d$, and $D=450d$, respectively. The drum is half-filled with $N$ particles, where $N$ are set to $7700$, $30600$, and $69200$ for $D=150d,300d$, and $450d$, respectively. Through this numerical simulation, we calculate the velocity distribution in the steady state using the procedure shown in Appendix A.
II.3 CFD simulation
In the CFD simulation, we treat the granular bed in the rotating drum as an incompressible non-Newtonian fluid. The continuum equations for the granular material are numerically solved using the finite difference method with the time step $\Delta t$ and grid width $\Delta x$. The free space in the drum is considered the gas phase to treat the free surface of the granular bed. The time evolution of the two-phase flow with the granular and gas phases is numerically solved by the volume of fluid (VoF) method [28, 29, 30]. This CFD simulation is an extension of the method implemented in Ref. [31] for an incompressible Newtonian fluid in a rotating container.
II.3.1 Governing equations
The entire domain $\mathrm{V}$ of the system is expressed as $\mathrm{V}=\mathrm{V_{f}}\cup\mathrm{V_{s}}$ with the fluid domain $\mathrm{V_{f}}$ including the granular and gas phases and the solid domain $\mathrm{V_{s}}$ representing the drum wall. The mass conservation equation is given by
$\nabla\cdot{\bf\it v}=0\;\;\;\;\;\mathrm{in}\;\;\;{\bf\it r}\in\mathrm{V},$ (19)
while the momentum conservation equation in the fluid domain is given by
$\frac{\partial{\bf\it v}}{\partial t}=-\nabla\cdot\left({\bf\it v}{\bf\it v}\right)-\frac{1}{\rho}\nabla p+\frac{1}{\rho}\nabla\cdot{\bf\it\tau}+{\bf\it g}\;\;\;\;\;\mathrm{in}\;\;\;{\bf\it r}\in\mathrm{V_{f}},$ (20)
where $\rho$, ${\bf\it v}$, and $p$ are the density, velocity, and pressure at ${\bf\it r}=(X,Z)$ and time $t$, respectively. Here, ${\bf\it\tau}$ is given by
${\bf\it\tau}=\eta{\bf\it\dot{\gamma}}$ (21)
with the viscosity $\eta$ and shear rate tensor ${\bf\it\dot{\gamma}}=\nabla{\bf\it v}+(\nabla{\bf\it v})^{T}$. Notably, $\eta$ depends on the position in the fluid domain. The velocity vector in the solid domain is given by
${\bf\it v}={\bf\it v}_{\mathrm{wall}}\;\;\;\;\;\mathrm{in}\;\;\;{\bf\it r}\in\mathrm{V_{s}},$ (22)
with
${\bf\it v}_{\mathrm{wall}}=-\Omega Z{\bf\it e}_{X}+\Omega X{\bf\it e}_{Z}.$ (23)
II.3.2 Volume of fluid method
In the VoF method, we introduce the phase volume fraction $\alpha({\bf\it r},t)$ of the fluid phase. Here, $\alpha=1$ and $\alpha=0$ correspond to the areas occupied only by the granular and gas phases, respectively. In the interface between the granular and gas phases, $\alpha$ satisfies $10<\alpha<$ 1. We consider the position with $\alpha=0.5$ as the free surface of the granular bed. The phase volume fraction $\alpha({\bf\it r},t)$ obeys the advection equation
$\frac{\partial\alpha}{\partial t}+\nabla\cdot\left(\alpha{\bf\it v}\right)-\alpha\nabla\cdot{\bf\it v}=0.$ (24)
This equation is solved using the WLIC/THINC (weighed line interface calculation/tangent of hyperbola for interface capturing) method [32]. The density $\rho$ and viscosity $\eta$ are calculated as follows:
$\rho$ $=\rho_{\mathrm{g}}\alpha+\rho_{\mathrm{air}}\left(1-\alpha\right),$ (25)
$\eta$ $=\eta_{\mathrm{g}}\left(\dot{\gamma},p\right)\alpha+\eta_{\mathrm{air}}\left(1-\alpha\right).$ (26)
Here, $\rho_{\mathrm{air}}$ and $\eta_{\mathrm{air}}$ are constants representing the density and viscosity in the gas phase, respectively. The density $\rho_{\mathrm{g}}$ in the granular phase is treated as a constant. The viscosity $\eta_{\mathrm{g}}$ in the granular phase is given by
$\eta_{\mathrm{g}}\left(\dot{\gamma},p\right)=\frac{\mu(I)p}{\dot{\gamma}}$ (27)
in the $\mu(I)$-rheology [19]. We adopt the functional form of $\mu(I)$ as Eq. (2). The inertial number $I$ is given by $I=\dot{\gamma}/\sqrt{p/\rho_{s}d^{2}}$ with
$\dot{\gamma}=\left(\frac{1}{2}{\bf\it\dot{\gamma}}:{\bf\it\dot{\gamma}}\right)^{\frac{1}{2}},$ (28)
which corresponds to the second invariant of ${\bf\it\dot{\gamma}}$. At constant pressure $p$, the $\mu(I)$-rheology corresponds to the Bingham model.
The viscosity in Eq. (27) diverges in the limit of $\dot{\gamma}\to 0$. To prevent the divergence in the CFD simulation, we regularize Eq. (27) by approximating it with a cutoff $\eta_{\mathrm{g,max}}$ [16] as follows:
$\eta_{\mathrm{g}}=\mathrm{min}\left[\frac{\mu(I)p}{\dot{\gamma}},\>\eta_{\mathrm{g,max}}\right].$ (29)
The rigid rotation in the static flow regime is effectively realized using a sufficiently large value for $\eta_{\mathrm{g,max}}$.
II.3.3 Simplified marker and cell method
In this study, we use the simplified marker and cell (SMAC) method [33] to solve the time evolution equations in the fluid domain. We denote an arbitrary physical quantity $\mathbf{a}$ at the time $t=n\Delta t$ by $\mathbf{a}^{n}$ with the time step $\Delta t$ and an integer $n$. In the SMAC method, $\mathbf{a}^{(n+1)}$ at the time $t=(n+1)\Delta t$ is numerically obtained from $\mathbf{a}^{(n)}$ in two steps.
First, for the prediction step, the predicted velocity $\mathbf{v}^{*}$ in ${\bf\it r}\in\mathrm{V_{f}}$ is determined by
$\frac{{\bf\it v}^{*}-{\bf\it v}^{n}}{\Delta t}=\frac{3{\bf\it A}^{n}-{\bf\it A}^{n-1}}{2}-\frac{1}{\rho^{n}}\nabla S^{n}+\frac{{\bf\it B}^{*}+{\bf\it B}^{n}}{2}+{\bf\it g},$ (30)
Here, we have used the second-order Adams–Bashforth method for the advection term
${\bf\it A}=-\nabla\cdot\left({\bf\it vv}\right),$ (31)
while the second-order Crank–Nicolson method is used for the viscous term
${\bf\it B}=\frac{1}{\rho}\nabla\cdot\left[\eta\left(\nabla{\bf\it v}+(\nabla{\bf\it v})^{T}\right)\right].$ (32)
In the SMAC method, a scholar potential $S$ is introduced in Eq. (30) to satisfy the mass conservation. In Eq. (30), ${\bf\it B}^{*}$ is given by
${\bf\it B}^{*}=\nu_{0}\nabla^{2}{\bf\it v}^{*}+\frac{1}{\rho^{n}}\nabla\cdot\left[\eta^{n}\left(\nabla{\bf\it v}^{n}+(\nabla{{\bf\it v}^{n}})^{T}\right)\right]-\nu_{0}\nabla^{2}{\bf\it v}^{n},$ (33)
where the first term is treated implicitly, and the other terms are treated explicitly [34]. Here, $\nu_{0}$ is the constant component of kinematic viscosity given by
$\nu_{0}=\frac{1}{2}\left(\frac{\eta_{\mathrm{air}}}{\rho_{\mathrm{air}}}+\frac{\eta_{\mathrm{g,max}}}{\rho_{\mathrm{g}}}\right).$ (34)
For ${\bf\it r}\in\mathrm{V_{s}}$, the predicted velocity $\mathbf{v}^{*}$ is given by
${\bf\it v}^{*}={\bf\it v}_{\mathrm{wall}}.$ (35)
Next, for the projection step, ${\bf\it v}^{\mathrm{n+1}}$ in ${\bf\it r}\in\mathrm{V_{f}}$ is obtained as follows:
${\bf\it v}^{\mathrm{n+1}}={\bf\it v}^{*}-\frac{\Delta t}{\rho^{n}}\nabla\sigma.$ (36)
Here, $\sigma$ is given by the solution of the Poisson equation
$\nabla\cdot\left(\frac{1}{\rho^{n}}\nabla\sigma\right)=\frac{1}{\Delta t}\nabla\cdot{\bf\it v}^{*}$ (37)
with the boundary condition $\nabla\sigma={\bf\it 0}$ at the wall surface, which is derived by taking the divergence of Eq. (36) with the mass conservation $\nabla\cdot{\bf\it v}^{n+1}=0$. The scholar potential $S^{n+1}$ is updated as follows:
$S^{n+1}=S^{n}+\sigma-\frac{1}{2}\Delta t\nu_{0}\nabla^{2}\sigma.$ (38)
In ${\bf\it r}\in\mathrm{V_{s}}$, ${\bf\it v}^{\mathrm{n+1}}$ is given by
${\bf\it v}^{\mathrm{n+1}}={\bf\it v}^{*}.$ (39)
According to Ref. [33], we calculate the pressure $p^{n}$ used in Eq. (29) as the solution of
$\nabla\cdot\left\{\frac{1}{\rho^{n}}\nabla p^{n}\right\}=\nabla\cdot\left(-\nabla\cdot\left({\bf\it v}^{n}{\bf\it v}^{n}\right)+\frac{1}{\rho^{n}}\nabla\cdot{\bf\it\tau}^{n}+{\bf\it g}\right)$ (40)
with the boundary conditions $\nabla p=\alpha\rho{\bf\it g}$ at the wall surface, which is derived from the divergence of the momentum conservation Eq. (20).
II.3.4 Boundary data immersion method
We treat the flow driven by the rotating drum using the boundary data immersion (BDI) method [35]. The solid domain $\mathrm{V_{s}}$ is defined as the circular shell with the inner and outer diameters $D$ and $D+16\Delta x$, respectively. In this method, we introduce a single meta-equation to solve the time evolution equations in the fluid and solid domains simultaneously. Here, we define the kernel function
$\chi(d_{k})=\begin{cases}1\;\;&\mathrm{for}\;\;d_{k}\leq-\epsilon\\ \frac{1}{2}\left[1-\sin\left(\frac{\pi d_{k}}{2\epsilon}\right)\right]\;\;&\mathrm{for}\;\;\left|d_{k}\right|<\epsilon\\ 0\;\;&d_{k}\geq\epsilon\end{cases}$ (41)
to continuously connect the fluid and solid domains. Here, the shortest distance from the wall surface at the position ${\bf\it r}$ is denoted by $d_{k}({\bf\it r})$, and $\epsilon$ is a parameter characterizing the thickness of the interface between the fluid and solid domains. From Eq. (41), the kernel function $\chi$ is $1$ in the fluid domain and 00 in the solid domain.
Using the BDI method with Eqs. (30) and (35), the meta-equation for the predicted velocity ${\bf\it v}^{*}$ is given by
${\bf\it v}^{*}=\chi\left[\left(1-\frac{\nu_{0}\Delta t}{2}\nabla^{2}\right)^{-1}{\bf\it\mathrm{RHS}}^{n}\right]+\left(1-\chi\right){\bf\it v}_{\mathrm{wall}},$ (42)
where
${\bf\it\mathrm{RHS}}^{n}={\bf\it v}^{n}+\Delta t\left[\frac{3{\bf\it A}^{n}-{\bf\it A}^{n-1}}{2}-\frac{1}{\rho^{n}}\nabla S^{n}+{\bf\it B}^{n}-\frac{\nu_{0}}{2}\nabla^{2}{\bf\it v}^{n}+{\bf\it g}\right].$ (43)
For the projection step, the meta-equation for ${\bf\it v}^{\mathrm{n+1}}$ is derived as follows:
${\bf\it v}^{\mathrm{n+1}}={\bf\it v}^{*}-\frac{\Delta t}{\rho^{n}}\chi\nabla\sigma$ (44)
from Eqs. (36) and (39). The Poisson equation for $\sigma$ is obtained as follows:
$\nabla\cdot\left(\frac{\chi}{\rho^{n}}\nabla\sigma\right)=\frac{1}{\Delta t}\nabla\cdot{\bf\it v}^{*},$ (45)
from the divergence of Eq. (44). The Poisson equation for $p^{n}$ is given by
$\nabla\cdot\left\{\frac{\chi}{\rho^{n}}\nabla p^{n}\right\}=\nabla\cdot\left\{\chi\left(-\nabla\cdot\left({\bf\it v}^{n}{\bf\it v}^{n}\right)+\frac{1}{\rho^{n}}\nabla\cdot{\bf\it\tau}^{n}+{\bf\it g}\right)+\left(1-\chi\right)\frac{{\bf\it v}_{\mathrm{wall}}}{\Delta t}\right\},$ (46)
which is obtained from the divergence of
${\bf\it v}^{n+1}=\chi\left[{\bf\it v}^{n}+\Delta t\left\{-\nabla\left({\bf\it v}^{n}{\bf\it v}^{n}\right)-\frac{1}{\rho^{n}}\nabla p^{n}+\frac{1}{\rho}\nabla\cdot{\bf\it\tau}^{n}+{\bf\it g}\right\}\right]+\left(1-\chi\right){\bf\it v}_{\mathrm{wall}}$ (47)
with Eqs. (20) and (22). We solve Eqs. (45) and (46) using the Bi-CGSTAB method [36].
Note that near the interface between the granular and gas phases, the solution of Eq. (46) may result in pressure below 00, which leads to unphysical behavior when calculating the viscosity of the granular phase $\eta_{\mathrm{g}}$ in Eq. (29) [37, 38]. To prevent this, a sufficiently small positive constant $p_{\mathrm{min}}$ is set as the minimum value of the pressure.
II.3.5 Numerical conditions
We estimate $\mu_{s}$, $\mu_{2}$, and $I_{0}$ in Eq. (2) as $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$ from the DEM simulation under uniform shear with the parameters in Sec. II.2.2. The details of the estimation are shown in Appendix B. Note that Eq. (2) is also used as the constitutive equation for grains with the Hertzian contact model by choosing appropriate values of the parameters [18]. We set the granular phase density as $\rho_{g}=\rho_{s}\phi$ with $\rho_{s}=m/(\pi d^{2}/4)$. The packing fraction $\phi$ of the particle is set to $0.825$ based on its value in the static flow regime of the rotating drum in the DEM simulation (Appendix C). We use $\eta_{\mathrm{g,max}}=8000m\sqrt{g/d}$, $\rho_{\mathrm{air}}=10^{-3}\rho_{g}$, $\eta_{\mathrm{air}}=10^{-6}\eta_{\mathrm{g,max}}$, and $p_{\mathrm{min}}=10^{-5}mg/d$. We have checked that the velocity profile is quantitatively the same even if we use larger $\eta_{\mathrm{g,max}}$ and lower $\rho_{\mathrm{air}}$, $\eta_{\mathrm{air}}$, and $p_{\mathrm{min}}$. We set $\Delta x=D/75$, $\Delta t=1.0\times 10^{-4}\sqrt{d/g}$, and $\epsilon=2\Delta x$. The maximum error for the Bi-CGSTAB method is set to $e_{\rm max}=10^{-5}$.
In this simulation, the free surface of the granular phase is initially set to be aligned with the $X$ axis, i.e., $\alpha=1$ for $Z<0$ and $\alpha=0$ for $Z\geq 0$. The initial velocity is 00 in the entire system. From this initial state, we generate a steady flow by rotating the drum counterclockwise with the angular velocity $\Omega$.
III Velocity profile
Figures 2 (a) and (b) show the velocity field of granular material for $D=150d$ and $\Omega=6.28\times 10^{-3}\sqrt{g/d}$ obtained from the DEM and CFD simulations, respectively. The red solid lines in these figures represent the free surface. In both cases, the granular material convects in a counterclockwise direction. Near the wall, the flow exhibits the rigid rotation following the rotating drum. Near the free surface, there is a region where the velocity is almost parallel to the free surface.
We set another Cartesian coordinate system $(x,z)$, where the $x$ axis is parallel to the free surface at $X=0$, as shown in Fig. 1. In this coordinate system, the depth from the free surface is represented by $z$. The coordinate system $(x,z)$ is given by
[…]
Here, $\theta$ denotes the angle of the free surface at $X=0$. The position of the free surface in the DEM simulation is determined from the packing fraction, as shown in Appendix C. Note that the free surface position at $x=0$ is treated as $z=0$, which is numerically evaluated in Appendix D.
Next, we analyze the velocity $u$ in the $x$-direction as a function of $z$ at the center of the granular bed, i.e. $x=0$. The velocity profile $u(z)$ for $D=150d$, $200d$, and $300d$ with $\Omega=6.28\times 10^{-3}\sqrt{g/d}$ is shown in Fig. 3 (a). The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The results of the CFD simulation are quantitatively consistent with those of the DEM simulation. The velocity $u(z)$ decreases as $z$ increases. For larger $z$, $u(z)$ coincides with the velocity $-\Omega z$ for the rigid rotation following the rotating drum, which is shown as the solid line in Fig. 3. This region is regarded as the static flow regime. Near $z=0$, the deviation from the rigid rotation becomes significant. This region corresponds to the surface flow layer. At the boundary between these two phases, $u(z)$ is close to 00, and the two phases are smoothly connected. These results are consistent with those in Ref. [10].
In Fig. 3 (b), the velocity profile $u(z)$ obtained from the DEM and CFD simulations for various values of $\Omega$ with $D=150d$ is shown. The results of the DEM and CFD simulations are also quantitatively consistent in this figure. For each $\Omega$, the surface flow layer and the static flow regime are distinguished. The magnitude and the slope of $u(z)$ in the surface flow layer increase with $\Omega$. Figure 3 demonstrates that the continuum model using the $\mu(I)$-rheology quantitatively reproduces the velocity field over a broad parameter space, extending beyond the limited conditions examined in previous studies [23, 13, 14, 16].
[…]
Figure 2: Velocity fields obtained from the DEM (a) and CFD (b) simulations with $D=150d$ and $\Omega=1.26\times 10^{-3}\sqrt{g/d}$. The red solid line represents the free surface.
[…]
Figure 3: (a) Velocity profile $u(z)$ at $x=0$ for $\Omega=6.28\times 10^{-3}\sqrt{g/d}$ with $D=150d$, $200d$, and $300d$. (b) Velocity profile $u(z)$ at $x=0$ for $D=150d$ with $\Omega=1.26\times 10^{-3}\sqrt{g/d}$ to $1.26\times 10^{-2}\sqrt{g/d}$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The solid lines represent $-\Omega z$ for the rigid rotation.
Next, to investigate the dependence of the velocity profile on $D$ and $\Omega$, we define the flow layer thickness $h$. Following Refs. [5, 7], we define the point with $u(z)=0$ in Fig. 3 as the boundary between the two phases and estimate the surface flow layer thickness $h$ as the distance from the free surface to the boundary point at $x=0$, which indicates
$u(z=h)=0.$ (49)
The surface flow layer thickness $h$ obtained from the DEM and CFD simulations is shown as a function of $\Omega$ for different $D$ in Fig. 4. The numerical results obtained from the DEM and CFD simulations show good agreement. The flow layer thickness $h$ increases with $D$ and depends slightly on $\Omega$. Note that the boundary has been estimated using different methods in previous studies [39, 40, 23, 10, 41]. However, the choice of the methods does not affect the qualitative behavior of $h$ as reported in Ref. [5]; this can be explained by the scaling of the velocity profile shown below (Appendix E).
[…]
Figure 4: Surface flow layer thickness $h$ against the angular velocity $\Omega$ for various values of $D$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively.
IV Scaling laws
In this section, we theoretically derive scaling laws for the velocity profile of granular flow in a rotating drum using the continuum model. We consider the mass conservation equation (19) and the momentum conservation equation (20). The viscosity $\eta$ in Eq. (20) is given by Eq. (27) with Eq. (2). We neglect the first term on the left-hand side of Eq. (20) by considering the steady flow. Using the drum diameter $D$, angular velocity $\Omega$, and density $\rho_{\mathrm{g}}$, we non-dimensionalize ${\bf\it r}$, $\nabla$, ${\bf\it v}$, $\dot{\gamma}$, $p$, and $\eta_{\mathrm{g}}$ as follows:
$\tilde{{\bf\it r}}=\frac{{\bf\it r}}{D},\;\;\tilde{\nabla}=\frac{\nabla}{D},\;\;\tilde{{\bf\it v}}=\frac{{\bf\it v}}{\Omega D},\;\;\tilde{\dot{\gamma}}=\frac{\dot{\gamma}}{\Omega},\;\;\tilde{p}=\frac{p}{\rho_{\mathrm{g}}\Omega^{2}D^{2}},\;\;\tilde{\eta}_{\mathrm{g}}=\frac{\eta_{\mathrm{g}}}{\rho_{\mathrm{g}}\Omega D^{2}}$ (50)
with the dimensionless variables $\tilde{{\bf\it r}},\tilde{\nabla},\tilde{{\bf\it v}},\tilde{\dot{\gamma}},\tilde{p}$, and $\tilde{\eta}_{\mathrm{g}}$. Here, we have used $D$ as a unit of length because we focus on the overall velocity profile in the system with $D\gg d$. Substituting these equations into Eqs. (19), (20), and (27) with Eq. (2), we obtain
~$\tilde{\nabla}\cdot\tilde{{\bf\it v}}=0,$ (51)
~$-\tilde{\nabla}\cdot\left(\tilde{{\bf\it v}}\tilde{{\bf\it v}}\right)-\tilde{\nabla}\tilde{p}+\tilde{\nabla}\cdot\left[\tilde{\eta_{\mathrm{g}}}\left(\tilde{\nabla}\tilde{{\bf\it v}}+\left(\tilde{\nabla}\tilde{{\bf\it v}}\right)^{T}\right)\right]-\frac{2}{Fr}{\bf\it e}_{z}=0,$ (52)
$\tilde{\eta}_{\mathrm{g}}=\mu_{s}\frac{\tilde{p}}{\tilde{\dot{\gamma}}}+\left(\mu_{2}-\mu_{s}\right)\frac{\tilde{p}}{I_{0}\sqrt{\phi\tilde{p}}/(d/D)+\tilde{\dot{\gamma}}},$ (53)
where we have introduced the unit vector ${\bf\it e}_{z}$ in the $z$ direction and the Froude number
$Fr=\frac{\Omega^{2}D}{2g}.$ (54)
Here, we consider the dependence of the velocity profile on $D$ and $\Omega$, and $d/D$ and $Fr$ are the dimensionless parameters related to them in Eqs. (51)-(53). Thus, the dimensionless velocity field $\tilde{{\bf\it v}}({\bf\it r})$ is characterized by the Froude number $Fr$ and the diameter ratio $d/D$, except for $\mu_{s}$, $\mu_{2}$, and $I_{0}$ determined by the properties of particles; this is consistent with the experiments in Ref. [42].
Since we have studied the system for $d/D\leq 1/150\ll 1$ in the previous section, we consider the limit $d/D\to 0$ in Eqs. (51)-(53). In this limit, the only dimensionless parameter in Eqs. (51)-(53) is $Fr$. Hence, the solution $\tilde{{\bf\it v}}\left(\tilde{{\bf\it r}}\right)$ of these equation depends only on $Fr$ and is written as
$\tilde{{\bf\it v}}\left(\tilde{{\bf\it r}}\right)={\bf\it V}\left(\tilde{{\bf\it r}};Fr\right),$ (55)
where ${\bf\it V}\left(\tilde{{\bf\it r}};Fr\right)$ is a function depending on $Fr$. Taking the limit $d/D\to 0$ in Eq. (53) is equivalent to treating the granular material as a simple plastic body with the yield stress proportional to the pressure. A similar approximation has been introduced in Ref. [43]. From Eq. (55), the scaling law for the velocity profile $u(z)$ at $x=0$ is obtained as follows:
$\frac{u(z)}{\Omega D}=U\left(\frac{z}{D};Fr\right),$ (56)
where $U\left(z/D;Fr\right)$ is a scaling function.
Substituting Eq. (56) into Eq. (49), we obtain
$U\left(\frac{h}{D};Fr\right)=0.$ (57)
From this equation, we derive the scaling law for $h$ as follows:
$\frac{h}{D}=H\left(Fr\right),$ (58)
where $H(Fr)$ is a scaling function.
We verify these scaling laws from the numerical simulations. Figure 5 illustrates the normalized velocity profile $u/\Omega D$ against $z/D$ for different $D/d$ with (a) $Fr=3.00\times 10^{-3}$ and (b) $Fr=1.20\times 10^{-2}$. The closed and open symbols represent the results of the CFD and DEM simulations, respectively. The solid line corresponds to the rigid rotation. We find excellent data consolidation in Fig. 5, which verifies the validity of the scaling law, Eq. (56). Comparing the results for different $Fr$, we find that the scaling function $U\left(z/D;Fr\right)$ shows a slight dependence on $Fr$.
[…]
Figure 5: Normalized velocity $u/\Omega D$ against $z/D$ at position $x=0$ for (a) $Fr=3.00\times 10^{-3}$, and (b) $Fr=1.20\times 10^{-2}$ with various values of $D$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The solid line corresponds to the rigid rotation.
Next, we plot $h/D$ against $Fr$ for different $d/D$ in Fig. 6. The closed and open symbols indicate the results of the CFD and DEM simulations, respectively. This figure verifies the validity of the scaling law, Eq. (58). We have focused on the steady flow for low $Fr<10^{-2}$. In this region, the scaling function $H(Fr)$ in Eq. (58) is expanded as follows:
$H\left(Fr\right)=H^{(0)}+FrH^{(1)}+O(Fr^{2})$ (59)
where $H^{(0)}$ and $H^{(1)}$ are constants. As shown in Appendix F, $H^{(0)}$ satisfies $H^{(0)}>0$ from the conservation of the flux in the steady flow. The scaling laws (58) with Eq. (59) derived from the continuum model explain that $h/D$ is almost constant with $Fr$ for $Fr\ll 1$ as shown in Fig. 6. This scaling law indicates that the surface layer thickness $h$ is proportional to $D$ but independent of $\Omega$ for sufficiently small $Fr$. This dependence of $h$ on $D$ and $\Omega$ is consistent with the previous experiment in Ref. [2].
[…]
Figure 6: Normalized surface flow layer thickness $h/D$ against the Froude number $Fr$ for various values of $D$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The solid line represents Eq. (59) with $H^{(0)}=1.1$ and $H^{(1)}=1.2$, which are estimated by the least squares method.
V Conclusion
In this paper, we numerically studied the steady flow of a two-dimensional granular material in a rotating drum using the DEM and CFD simulations with the $\mu(I)$-rheology. The velocity fields obtained from both methods are quantitatively consistent for various values of the diameter $D$ and the angular velocity $\Omega$ of the drum. The granular material exhibits the rigid rotation in the static flow regime near the drum wall while it flows parallel to the free surface in the surface flow layer. The surface flow layer is characterized by the layer thickness $h$ at the center of the drum. We have analytically obtained scaling laws (Eqs. (56) and (58)) for the velocity profile and surface layer thickness from dimensional considerations of the continuum equations. Numerical results verified the validity of these scaling laws.
The scaling law for the surface flow layer thickness based on the $\mu(I)$-rheology predicts $h\sim\mbox{const.}$. in the limit of $\Omega\to 0$, which is consistent with the experimental results in Ref. [2]. In contrast, other experimental studies [3, 5, 6, 7] have reported a power-law dependence, $h\sim\Omega^{\alpha}$ with a constant $\alpha$, which was obtained from a phenomenological assumption of a constant shear rate. The cylinder used in Ref. [2] is sufficiently long in the axial direction, and the flow can be considered two-dimensional, which may explain the agreement with our theoretical prediction. Our preliminary three-dimensional simulations suggest that the scaling law (Eqs. (56) and (58)) holds for granular flow in a sufficiently long rotating cylinder, whereas significant deviations are observed for shorter cylinders. These deviations may be attributed to the influence of sidewall friction, potentially explaining why the experimental results in Refs. [3, 5, 6, 7] deviate from our scaling laws. However, the effect of sidewall friction in shorter cylinders and a theoretical explanation for the experimental results in Refs. [3, 5, 6, 7] warrant further careful investigation.
We studied the rotating drum half-filled with granular materials in this study. However, a flow similar to Fig. 1 is observed for various values of the fill fraction. We have not investigated the dependence of the velocity field on the fill fraction; however, we expect the scaling laws derived in this study to apply to systems with different fill fractions because we have not used the condition for the fill fractions in the derivation of the scaling law for the velocity field ${\bf\it v}({\bf\it r})$.
Recently, the $\mu(I)$-rheology has been improved to include the non-local effect [20, 22]. We ignored its effect and adopted the local $\mu(I)$-rheology in our analysis since we deemed it irrelevant for sufficiently large systems with $D/d\geq 150$. In fact, the $\mu(I)$-rheology has qualitatively reproduced the velocity fields in our simulation. However, more detailed work is necessary on the non-local effect on the velocity field in the rotating drum.
Although we obtained the scaling laws by the non-dimensionalization, the forms of the scaling functions are not theoretically derived. In Ref. [44], the velocity field of an incompressible Newtonian fluid in a rotating drum is analytically calculated. In. Ref. [45], the streamlines for the granular flow in a rotating drum are derived based on a simple constitutive law. The development of these analytical methods may lead to a better understanding of the granular flow, including the derivation of the scaling functions.
Acknowledgements.
The authors thank S. Goto, Y. Motoori, D. Watanabe, and I. Procaccia for fruitful discussions. We would like to thank Editage (www.editage.jp) for English language editing. This study was supported by JSPS KAKENHI under Grant No. JP23K03248 and JP21H01006. H.O. conducted the simulations and analysis. All authors analyzed the results and reviewed the manuscript.
Appendix A Velocity field in DEM simulation
The velocity distribution in Fig. 2(a) is calculated using the following procedure in the DEM simulation. First, the system is divided into small square regions with a length of $4.8d$ in the plane of the Cartesian coordinates $\left(X,Z\right)$. Then, the average particle velocity in each region is plotted as the velocity field.
In the DEM simulation, the velocity field in the plane of the Cartesian coordinates $(x,z)$ shown in Fig. 3 is numerically obtained as the average particle velocity in the square regions with a length of $1.2d$. In the CFD simulation, the velocity at each region is calculated as the average of the values on the computational grids overlapping with the region.
Appendix B Determination of parameters for $\mu(I)$-rheology
To estimate the parameters in Eq. (2), we have performed the DEM simulation for the granular materials shown in Sec. II.2 under uniform shear with the shear rate $\dot{\gamma}$ and packing fraction $\phi$ using the Lees-Edwards boundary condition [46]. We use $\phi=0.825$, which is the value in the static flow regime of the rotating drum. In this simulation, the shear stress $\tau$ and pressure $p$ are measured, and we plot the bulk friction $\mu=\tau/p$ against the inertial number $I$ in Fig. 7. This figure shows that the $I$-dependence of $\mu$ is consistent with the theoretical expression, Eq. (2), for the $\mu(I)$-rheology represented by the solid line. We have determined the parameters as $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$ by fitting the numerical data with Eq. (2).
[…]
Figure 7: Bulk friction $\mu$ against the inertial number $I$ obtained from the DEM simulation under uniform shear. The solid line represents Eq. (2) with $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$.
Appendix C Packing fraction in the rotating drum
Figure 8 presents the dependence of the volume fraction $\phi$ on $z$ at $x=0$ obtained from the DEM simulation with $D=150d$ and $\Omega=6.28\times 10^{-3}\sqrt{\rho_{g}g/d}$. The packing fraction $\phi$ is nearly constant except for the region near $z\approx 0$. The packing fraction in the static flow regime is estimated to be $0.825$, which is indicated by the straight line in Fig. 8, though the packing fraction in the surface flow layer is slightly small due to the dilatancy of granular materials under shear [10]. Therefore, in the CFD simulation, the granular material is treated as an incompressible fluid. In the DEM simulation, $\phi=0.7$ is defined as the free surface position.
[…]
Figure 8: Packing fraction $\phi$ against $z$ at $x=0$ obtained from the DEM simulation with $D=150d$ and $\Omega=6.28\times 10^{-3}\sqrt{\rho_{g}g/d}$. The solid line represents the average value of $\phi$ in the solid region.
Appendix D Position of free surface
Figure 9 illustrates the position of the free surface $z_{\mathrm{surface}}$ against the Froude number $Fr$ at $x=0$ for each $D$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The normalized free surface position $z_{\mathrm{surface}}/D$ is close to 00 for all of the Froude number $Fr$. Therefore, we treat $z_{\mathrm{surface}}=0$ in this study.
[…]
Figure 9: Normalized free surface height $z_{\mathrm{surface}}/D$ at $x=0$ against the Froude number $Fr$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively.
Appendix E Boundary between surface flow layer and static flow regime
In this study, the boundary between the surface flow layer and the static flow regime is defined as the position for $u(z)=0$. However, in Ref. [39], the boundary is determined from the difference between $u(z)$ and the rigid rotation velocity $-\Omega z$. For example, we may regard the surface flow layer as the region where the velocity in the granular bed is sufficiently larger than that of the rigid rotation as
$\frac{u(z)-\left(-\Omega z\right)}{\Omega D}\geq\alpha_{\mathrm{th}}$ (60)
with a threshold $\alpha_{\mathrm{th}}$. The surface flow layer thickness $h^{\prime}$ for this definition satisfies
$\frac{u(z=h^{\prime})-\left(-\Omega h^{\prime}\right)}{\Omega D}=\alpha_{\mathrm{th}}.$ (61)
If the velocity $u(z)$ obeys the scaling law, Eq. (56), this equation gives
$U\left(\frac{h^{\prime}}{D};Fr\right)-\frac{h^{\prime}}{D}=\alpha_{\mathrm{th}}.$ (62)
Solving this equation for $h^{\prime}/D$ yields the scaling law for $h^{\prime}$ as
$\frac{h^{\prime}}{D}=H^{\prime}\left(Fr\right)$ (63)
for a given $\alpha_{\mathrm{th}}$, where $H^{\prime}(Fr)$ is a scaling function. This equation is qualitatively consistent with the scaling law (58) for $h$. This indicates that the surface flow layer thickness shows the same scaling regardless of the definition of $h$.
Appendix F Surface flow layer thickness $h$ for $Fr\ll 1$
In this section, we show $H^{(0)}>0$ in Eq. (59). The velocity $u(z)$ is a decreasing function of $z$ and $u(z)=0$ at $z=h$. The volumetric flow rates in the surface flow layer and the static flow regime are given by
[…]
and
[…]
respectively. Since the system is in a steady state, these volumetric flow rates satisfy
$Q_{+}=Q_{-}.$ (66)
The velocity $u(z)$ satisfies Eq. (56) and its scaling function $U(\tilde{z}:Fr)$ is a decreasing function of $\tilde{z}=z/D$ with $U(\tilde{z}=H(Fr);Fr)=0$. Therefore, $U(\tilde{z}:Fr)$ is represented as
$U(\tilde{z}:Fr)=\left\{\begin{array}[]{ll}U_{+}(H(Fr)-\tilde{z};Fr)&\tilde{z}<H(Fr)\\ -U_{-}(\tilde{z}-H(Fr);Fr)&\tilde{z}\geq H(Fr)\end{array}\right.,$ (69)
where $U_{\pm}(\zeta;Fr)$ is an increasing function of $\zeta$ satisfying
$U_{\pm}(\zeta=0;Fr)=0,\quad U_{\pm}(\zeta>0;Fr)>0.$ (70)
Substituting Eqs. (56) and (69) into Eqs. (64) and (65), $Q_{\pm}$ is expressed as
[…]
Since $Fr\ll 1$, $U_{\pm}(\zeta;Fr)$ is expanded as follows:
$\begin{array}[]{ll}U_{\pm}(\zeta;Fr)=U_{\pm}^{(0)}(\zeta)+FrU_{\pm}^{(1)}(\zeta)+O(Fr^{2}),\end{array}$ (73)
where $U_{\pm}^{(0)}(\zeta)$ and $U_{\pm}^{(1)}(\zeta)$ are functions depending only on $\zeta$. Considering $U_{\pm}(\zeta;Fr)$ in the limit of $Fr\to 0$ and Eq. (70), we find that $U_{\pm}^{(0)}(\zeta)$ satisfies
$U_{\pm}^{(0)}(\zeta=0)=0,\quad U_{\pm}^{(0)}(\zeta>0)>0.$ (74)
Substituting Eqs. (59), (71), (72), and (73) into Eq. (66), we obtain
[…]
Since this equation holds for any $Fr$, we obtain
[…]
Since $20\leq h/D\leq 1/$ 2 from the definition, $H^{(0)}$ satisfies $20\leq H^{(0)}\leq 1/$ 2. Here, $H^{(0)}$ is the solution of Eq. (76); however, $H^{(0)}=0$ does not satisfy this equation due to Eq. (70). Therefore, we obtain
$H^{(0)}>0.$ (77)
This equation with Eqs(59) and (58) explains the behavior of $h$ in Fig. 6.